### v1

@@ Abstract
The robustness of driving perception systems under unprecedented conditions is crucial for safety-critical usages. Latest advancements have prompted increasing interests towards multi-LiDAR perception. However, prevailing driving datasets predominantly utilize single-LiDAR systems and collect data devoid of adverse conditions, failing to capture the complexities of real-world environments accurately. Addressing these gaps, we proposed Place3D, a full-cycle pipeline that encompasses LiDAR placement optimization, data generation, and downstream evaluations. Our framework makes three appealing contributions. 1) To identify the most effective configurations for multi-LiDAR systems, we introduce a Surrogate Metric of the Semantic Occupancy Grids (M-SOG) to evaluate LiDAR placement quality. 2) Leveraging the M-SOG metric, we propose a novel optimization strategy to refine multi-LiDAR placements. 3) Centered around the theme of multi-condition multi-LiDAR perception, we collect a 364,000-frame dataset from both clean and adverse conditions. Extensive experiments demonstrate that LiDAR placements optimized using our approach outperform various baselines. We showcase exceptional robustness in both 3D object detection and LiDAR semantic segmentation tasks, under diverse adverse weather and sensor failure conditions. Code and benchmark toolkit are publicly available. [^1] GitHub Repo: https://github.com/ywyeli/Place3D.
Keywords:
Autonomous Driving Sensor Placement 3D Robustness
@@ 1 Introduction
(a) Waymo
(b) Motional
(c) Cruise
(d) Pony.ai
(e) Zoox
(f) Toyota
(g) Momenta
(h) Ford
Figure 1: A diverse spectrum of existing multi-LiDAR placement configurations employed by major autonomous vehicle companies [5, 4, 3, 2, 7, 8, 6, 1]. Images adopted from original websites.
Accurate 3D perception plays a crucial role in autonomous driving, involving detecting the objects around the vehicle and segmenting the scene into meaningful semantic categories. LiDARs are becoming crucial for driving perception due to their capability to capture detailed geometric information about the surroundings [11, 97]. While the latest models achieved promising accuracy on standard datasets, e.g., nuScenes [13], improving the robustness of perception under corruptions and sensor failures is still a critical yet under-explored task [46, 84].
Recent studies have primarily focused on refining the sensing systems by designing new algorithms with novel model architectures [19] or 3D representations [28, 88]. The selection of LiDAR configurations, however, often relies on industry experience and design aesthetics. Therefore, existing literature has potentially overlooked optimal LiDAR placements for maximum sensing efficacy.
An intuitive approach for optimizing LiDAR configurations involves a comprehensive cycle of data collection, model training, and validation across various LiDAR setups to enhance autonomous driving systems’ perception accuracy. However, this approach faces significant challenges due to the substantial computational resources and extensive time required for data collection and processing [53]. Although recent works [34, 60] made preliminary attempts to explore the impact of LiDAR placements on perception accuracy, they have neither proposed an optimization method nor evaluated the performance in adverse conditions.
@@
We first introduce an easy-to-compute Surrogate Metric of Semantic Occupancy Grids (M-SOG) to evaluate the equality of LiDAR placements. Next, we propose a novel optimization approach utilizing our surrogate metric based on Covariance Matrix Adaptation Evolution Strategy (CMA-ES) to find the near-optimal LiDAR placements. To verify the correlation between our surrogate metric and assess the effectiveness of our optimization approach on both clean and adverse conditions, we design an automated multi-condition multi-LiDAR data simulation platform and establish a comprehensive benchmark consisting of a total of seven LiDAR placement baselines inspired by existing self-driving configuration from the autonomous vehicle companies (cf. Fig. 1).
Our benchmark, along with a large-scale multi-condition multi-LiDAR perception dataset, encompasses state-of-the-art learning-based perception models for 3D object detection [50, 95, 58, 97] and LiDAR semantic segmentation [21, 99, 78, 105], as well as six distinct adverse conditions coped with weather and sensor failures. Utilizing the proposed Place3D framework, we explored how various perturbations and downstream 3D perception tasks affect optimization outcomes. To summarize, this work makes the following key contributions:
  • To the best of our knowledge, Place3D serves as the first attempt at investigating the impact of multi-LiDAR placements in diverse conditions.
  • We introduce M-SOG, an innovative surrogate metric to effectively evaluate the quality of LiDAR placements for both detection and segmentation tasks.
  • We propose a novel optimization approach utilizing our surrogate metric to refine LiDAR placements, which exhibit excellent 3D perception accuracy and robustness, outperforming baselines for up to 9% scores.
  • We contribute a 364,000-frame multi-condition multi-LiDAR point cloud dataset and establish a comprehensive benchmark for the 3D robustness evaluation. We hope this work can lay a solid foundation for future research.
@@ 2 Related Work
LiDAR Sensing. LiDAR sensing plays a vital role in advanced autonomous vehicles, acquiring structural information from the surroundings and serving as the most critical components for their operation [55, 11, 66]. Leveraging its 3D data, LiDAR enables a variety of tasks designed to meet practical needs. Prominent among these tasks include perception [13, 27, 29, 12, 77, 48], generation [68, 106, 67, 86], decision making [26, 14, 22, 79], and simulation [25, 63, 62, 83]. This work specifically targets LiDAR-based perception and simulation, which are at the forefront of current research in autonomous vehicle technology.
LiDAR Semantic Segmentation. As one of the most crucial tasks in LiDAR-based perception, LiDAR segmentation has attracted wide research attention. Based on representations of LiDAR points, the LiDAR segmentation models can be categorized into point-based [80, 35, 36, 98, 69], range view [65, 81, 24, 87, 101, 20, 47, 9, 45, 88], bird’s eye view [99, 103, 16], voxel-based [21, 78, 105, 33], and multi-view fusion [54, 38, 70, 56, 57, 18, 17] methods. Albeit achieving promising results on standard LiDAR semantic segmentation benchmarks, their performance across various LiDAR sensor configurations has not been thoroughly explored. In this work, we conduct the first study to examine how different LiDAR placements affect the effectiveness of state-of-the-art LiDAR segmentation algorithms.
3D Object Detection. Another critical task for driving perception aims to locate and classify road participants in 3D scenes [10]. Most prevailing 3D object detection models adopt point-based [74, 94, 93, 104] or voxel-based [102, 50, 64, 82, 75, 51, 61, 92] representations. Recently, there has been a shift towards fusing both approaches for improved detection performance [59, 72, 73, 58, 52]. Our study is perpendicular to existing 3D object detection approaches. We find that strategically optimized LiDAR placements can significantly enhance 3D object detection performance under a variety of real-world sensing conditions.
@@
Perception Robustness. The reliability of perception models is crucial for driving applications in real-world scenarios [76, 39]. Recently, several studies explored the robustness of 3D perception models against adversarial attacks [100, 85], common corruptions [46, 49, 84, 91, 43], adverse weather conditions [31, 30, 71, 37], sensor failures [96, 19, 28], and combined motion and sensor perturbations [90]. To the best of our knowledge, no prior work has specifically targeted optimizing the placement of LiDAR sensors to improve 3D robustness. Our Place3D fills this gap by introducing a comprehensive study of different placement strategies against both in-domain and out-of-domain scenarios, which could inspire future exploration in optimizing sensor configurations to fortify perception robustness.
@@
Sensor Placement Optimization. Optimizing sensor placements can alleviate the challenges of heuristic design and significantly enhance performance [42, 89]. In the realm of autonomous driving, optimizing LiDAR sensor placements represents a relatively new field study[60]. Hu et al. [34] presented the first study on optimizing multi-LiDAR placements for 3D object detection. Li et al. [53] focused on the influence of LiDAR-camera configurations on multi-modal 3D object detection. Several other works [41, 44, 15, 40] explored the placements of roadside LiDAR sensors for vehicle-to-everything applications, diverging from the focus on in-vehicle perception. Distinguishing from previous efforts, our work makes the first attempt at investigating LiDAR placements for robust 3D perception in challenging conditions. We establish comprehensive benchmarks for both LiDAR semantic segmentation and 3D object detection and perform in-depth analysis on the optimization of LiDAR placements for robust perception.
3 Place3D
In this section, we introduce the Surrogate Metric of Semantic Occupancy Grids (M-SOG) to assess the 3D perception performance of sensor configurations (cf. Sec. 3.1). Leveraging M-SOG, we propose a novel optimization strategy for multi-LiDAR placement (cf. Sec. 3.2). Our approach is theoretically grounded; we provide an optimality certification to ensure that optimized placements are close to the global optimum (cf. Sec. 3.3). Fig. 2 depicts an overview of our framework.
@@ 3 Place3D
Figure 2: Place3D pipeline for multi-LiDAR placement optimization. We first generate Probabilistic SOG (b) by merging multiple frames of data synthesized in CARLA into dense point cloud (a) and obtain voxels covered by LiDAR rays to compute the Surrogate Metric of Semantic Occupancy Grids (M-SOG) based on the joint probability distribution of covered voxels (c). We propose a CMA-ES-based optimization strategy to maximize M-SOG, finding optimal LiDAR placement (d). To verify the effectiveness of our optimization strategy, we contribute a multi-condition multi-LiDAR dataset (e). We train several state-of-the-art models and evaluate the performance of baseline and optimized placements on both clean and corruption data (f). Our optimized placement generally outperforms all baseline placements in experiments.
@@ 3.1 M-SOG: Surrogate Metric of Semantic Occupancy Grids
Region of Interest. We first define the Region of Interest (ROI) for perception as a finite 3D cuboid space $[L,W,H]$ with ego-vehicle in the center and divide the ROI space $S$ into voxels with resolution [$\delta_{L},\delta_{W},\delta_{H}$] as $S=\{v_{1},v_{2},...,v_{N}\}$, where $N=\frac{L}{\delta_{L}}\times\frac{W}{\delta_{W}}\times\frac{H}{\delta_{H}}$ denotes the total number of voxelized grids.
@@
Semantic Occupancy Grids (SOG). Given a 3D scene with a set of semantic labels $Y=\{y_{1},y_{2},...,y_{K}\}$, where $K$ represents the total number of object classes in the scene and assuming that each voxel can only be occupied by one semantic tag for each frame, we denote $y^{(t)}(v_{i})$ as the semantic label of voxel $v_{i}\in\{v_{1},v_{2},...,v_{N}\}$ at time frame $t\in\{1,2,...,T\}$. Accordingly, the set of voxels occupied by semantic label $y_{k}$ at frame $t$ is defined as $s_{y_{k}}^{(t)}=\{v_{i}|y^{(t)}(v_{i})=y_{k}\},\;k=1,2,...,K,\;t=1,2,...,T$. Then, we introduce the Semantic Occupancy Grids to describe the total semantic voxel distribution,
$S_{SOG}^{(t)}=\{s_{y_{1}}^{(t)},s_{y_{2}}^{(t)},...,s_{y_{K}}^{(t)}\},\;t=1,2,...,T~{}.$ (1)
@@
Probabilistic SOG. We propose P-SOG $p_{SOG|y_{k}}=p\{v_{1},v_{2},...,v_{N}|y_{k}\}$ to describe the joint probability distribution of voxels belonging to certain semantic labels. Before estimating $p_{SOG|y_{k}}$, we first traverse all frames from given scenes to obtain the probability $\hat{p}$ for each voxel $v_{i}$ occupied by each semantic label:
$\hat{p}(v_{i}|y_{k})=\frac{\sum_{t=1}^{T}\mathds{1}(v_{i}\in s_{y_{k}}^{(t)})}{T}~{},$ (2)
@@
where $\mathds{1}(\cdot)$ serves as an occupancy indicator function. Notably, $\hat{p}$ indicates estimated distributions from observed samples, whereas $p$ refers to the statistical parameters to be estimated, which are unknown and non-random. We compute the joint probability of all non-zero voxels for semantic label $y_{k}$ in the ROI to estimate the $p_{SOG|y_{k}}$. Recognizing that an object’s presence in one voxel does not imply its presence in others, we treat voxels as independent, identically distributed variables, calculating the joint distribution over the ROI set $S$ as:
$\hat{p}_{SOG|y_{k}}=\hat{p}(v_{1},v_{2},...,v_{N}|y_{k})=\prod_{i=1,\;\hat{p}(v_{i}|y_{k})\neq 0}^{N}{\hat{p}(v_{i}|y_{k})}~{}.$ (3)
@@
Conditional P-SOG. To evaluate LiDAR placements, we further analyze the joint probability distribution of voxels covered by LiDAR rays. Leveraging the Bresenham’s Line Algorithm, we obtain the voxels covered by LiDAR rays given LiDAR configuration $L=L_{j}$, $j=1,2,...,J$ as $S|L_{j}=Bresenham(S,L_{j})=\{v_{1}^{L_{j}},v_{2}^{L_{j}},...,v^{L_{j}}_{N_{j}}\}$, where $J$ is the number of total potential configurations and $N_{j}$ is the number of voxels covered by rays of LiDAR Configuration $L=L_{j}$. Then we have the conditional P-SOG $\hat{p}_{SOG|y_{k},L_{j}}=\hat{p}(v_{1}^{L_{j}},v_{2}^{L_{j}},...,v_{N}^{L_{j}}|y_{k})$. From the perspective of density estimation to find the P-SOG of the semantic object $y_{k}$, the P-SOG and conditional P-SOG can be estimated as $p_{SOG|y_{k}}=\hat{p}_{SOG|y_{k}},p_{SOG|y_{k},L_{j}}=\hat{p}_{SOG|y_{k},L_{j}}$, respectively.
Figure 3: Pipeline of probabilistic SOG generation. We generate dense point clouds (b) by merging multiple frames of raw point clouds (a). Then, we voxelize the dense point clouds into Semantic Occupancy Grids (SOG) (c), and traverse all frames of dense point cloud to compute Probabilistic Semantic Occupancy Grids (P-SOG) (d).
Surrogate Metric of SOG (M-SOG). Entropy can be used to describe the uncertainty, indicating the perception capability of sensors. To this end, we calculate the entropy for each voxel, i.e., $\hat{H}(v_{i})=-\sum_{k=1}^{K}\hat{p}(v_{i}|y_{k})\log\hat{p}(v_{i}|y_{k})$. Then, we obtain the total entropy of P-SOG and conditional P-SOG as follows:
$H_{SOG}=H(S)$ $=\mathds{E}_{v_{i}\sim p_{S}}\sum_{i=1}^{N}\hat{H}(v_{i})~{},$ (4)
$H_{SOG|L_{j}}=H(S|L_{j})$ $=\mathds{E}_{v_{i}^{L_{j}}\sim p_{S|L_{j}}}\sum_{i=1}^{N_{j}}\hat{H}(v_{i}^{L_{j}})~{}.$ (5)
@@
We further define the Information Gain $IG_{S|L_{j}}=\frac{H(S)}{N}-\frac{H(S|L_{j})}{N_{j}}$ given LiDAR configuration $L_{j}$ to describe the difference between the mean values of entropy of P-SOG and conditional P-SOG. Since $\frac{H(S)}{N}$ is a constant given certain scenes with fixed P-SOG distribution, we propose the Surrogate Metric of Semantic Occupancy Grids (M-SOG) as follows to evaluate the perception capability:
$\mathbf{M}_{SOG}(L_{j})=-\frac{H(S|L_{j})}{N_{j}}~{}.$ (6)

### v2

@@ Abstract
The reliability of driving perception systems under unprecedented conditions is crucial for practical usage. Latest advancements have prompted increasing interest in multi-LiDAR perception. However, prevailing driving datasets predominantly utilize single-LiDAR systems and collect data devoid of adverse conditions, failing to capture the complexities of real-world environments accurately. Addressing these gaps, we proposed Place3D, a full-cycle pipeline that encompasses LiDAR placement optimization, data generation, and downstream evaluations. Our framework makes three appealing contributions. 1) To identify the most effective configurations for multi-LiDAR systems, we introduce the Surrogate Metric of the Semantic Occupancy Grids (M-SOG) to evaluate LiDAR placement quality. 2) Leveraging the M-SOG metric, we propose a novel optimization strategy to refine multi-LiDAR placements. 3) Centered around the theme of multi-condition multi-LiDAR perception, we collect a 280,000-frame dataset from both clean and adverse conditions. Extensive experiments demonstrate that LiDAR placements optimized using our approach outperform various baselines. We showcase exceptional results in both LiDAR semantic segmentation and 3D object detection tasks, under diverse weather and sensor failure conditions.
@@ 1 Introduction
Accurate 3D perception plays a crucial role in autonomous driving, involving detecting the objects around the vehicle and segmenting the scene into meaningful semantic categories. LiDARs are becoming crucial for driving perception due to their capability to capture detailed geometric information about the surroundings [3, 109, 53, 54, 118]. While the latest models achieved promising accuracy on standard datasets, e.g., nuScenes [7] and SemanticKITTI [6], improving the resilience of perception under corruptions and sensor failures is still a critical yet under-explored task [46, 94, 5, 95].
Recent studies have primarily focused on refining the sensing systems by designing new algorithms with novel model architectures [15] or 3D representations [28, 99, 11]. The selection of LiDAR configurations, however, often relies on industry experience and design aesthetics. Therefore, existing literature has potentially overlooked optimal LiDAR placements for maximum sensing efficacy.
An intuitive approach for optimizing LiDAR configurations involves a comprehensive cycle of data collection, model training, and validation across various LiDAR setups to enhance the autonomous driving system’s perception accuracy [24, 52]. However, this approach faces significant challenges due to the substantial computational resources and extensive time required for data collection and processing [57, 17]. Although recent works [34, 65] made preliminary attempts to explore the impact of LiDAR placements on perception accuracy, they have neither proposed an optimization method nor evaluated the performance in adverse conditions.
@@
We first introduce an easy-to-compute Surrogate Metric of Semantic Occupancy Grids (M-SOG) to evaluate the equality of LiDAR placements. Next, we propose a novel optimization approach utilizing our surrogate metric based on Covariance Matrix Adaptation Evolution Strategy (CMA-ES) to find the near-optimal LiDAR placements. To verify the correlation between our surrogate metric and assess the effectiveness of our optimization approach on both clean and adverse conditions, we design an automated multi-condition multi-LiDAR data simulation platform and establish a comprehensive benchmark consisting of a total of seven LiDAR placement baselines inspired by existing self-driving configuration from the autonomous vehicle companies.
Our benchmark, along with a large-scale multi-condition multi-LiDAR perception dataset, encompasses state-of-the-art learning-based perception models for 3D object detection [51, 107, 63, 109] and LiDAR semantic segmentation [18, 111, 86, 117, 50], as well as six distinct adverse conditions coped with weather and sensor failures [46, 95]. Utilizing the proposed framework, we explored how various perturbations and downstream 3D perception tasks affect optimization outcomes.
To summarize, this work makes the following key contributions:
  • To the best of our knowledge, Place3D serves as the first attempt at investigating the impact of multi-LiDAR placements for 3D semantic scene understanding in diverse conditions.
  • We introduce M-SOG, an innovative surrogate metric to effectively evaluate the quality of LiDAR placements for both detection (sparse 3D) and segmentation (dense 3D) tasks.
  • We propose a novel optimization approach utilizing our surrogate metric to refine LiDAR placements, which exhibit excellent LiDAR semantic segmentation and 3D object detection accuracy and robustness, outperforming baselines for relatively 9% in metrics.
  • We contribute a 280,000-frame multi-condition multi-LiDAR point cloud dataset and establish a comprehensive benchmark for LiDAR-based 3D scene understanding evaluations. We hope this work can lay a solid foundation for future research in this relevant field.
@@ 2 Related Work
LiDAR Sensing. LiDAR sensing is critical in autonomous vehicles, providing essential structural information for their operation [59, 3, 73]. Utilizing 3D data, LiDAR supports tasks such as semantic scene understanding [7, 26, 29, 6, 85, 48, 10], generation [75, 120, 74, 97], decision making [25, 8, 19, 87], and simulation [23, 68, 67, 93]. This work specifically targets LiDAR-based perception and simulation, which are at the forefront of current research in autonomous vehicle technology.
LiDAR-Based 3D Scene Understanding. LiDAR segmentation and 3D object detection are key tasks for 3D scene understanding. Segmentation models are categorized into point-based [88, 35, 36, 110, 77], range view [70, 90, 21, 98, 113, 16, 47, 2, 45, 99], bird’s eye view [111, 115, 12], voxel-based [18, 86, 117, 33], and multi-view fusion [58, 38, 78, 60, 62, 14, 13, 61, 100] methods. While they show promising results on benchmarks, their performance across different LiDAR configurations is not well explored. For 3D object detection, models typically use point-based [82, 106, 105, 116] or voxel-based [114, 51, 69, 92, 83, 55, 66, 104] representations, with recent works favoring fusion for improved detection [64, 80, 81, 63, 56]. Our study complements these approaches by optimizing LiDAR placements to enhance performance under various real-world conditions.
@@
LiDAR Perception Robustness. The reliability of LiDAR-based 3D scene understanding models in real-world scenarios is crucial [84, 39]. Recent studies have explored the robustness of 3D perception models against adversarial attacks [112, 96], common corruptions [46, 49, 94, 103, 43, 4], adverse weather [31, 30, 79, 37], sensor failures [108, 15, 28], and combined motion and sensor perturbations [102, 95]. To our knowledge, no prior work has focused on optimizing LiDAR placement to improve semantic 3D scene understanding robustness. Our Place3D addresses this gap by studying various placement strategies to enhance robustness in both in-domain and out-of-domain scenarios.
@@
Sensor Placement Optimization. Optimizing sensor placements can enhance performance and address the challenges of heuristic design [42, 101]. In autonomous driving, optimizing LiDAR placements is relatively new [65]. Hu et al. [34] studied multi-LiDAR placements for 3D object detection, while Li et al. [57] examined LiDAR-camera configurations for multi-modal detection. Other works [41, 44, 9, 40] explored roadside LiDAR placements for V2X applications. Distinguishing from these efforts, our work is the first to investigate LiDAR placements for robust 3D scene understanding in challenging conditions. We establish benchmarks for both LiDAR semantic segmentation and 3D object detection, providing an in-depth analysis of optimal placements for robust perception.
@@
Figure 1: Place3D pipeline for multi-LiDAR placement optimization. We first utilize the semantic point cloud synthesized in CARLA (a) to generate Probabilistic SOG (b) and obtain voxels covered by LiDAR rays to compute M-SOG (c). We propose a CMA-ES-based optimization strategy to maximize M-SOG, finding optimal LiDAR placement (d). To verify the effectiveness of our LiDAR placement optimization strategy, we contribute a multi-condition multi-LiDAR dataset (e) and evaluate the performance of baselines and optimized placements on both clean and corruption data (f).
3 Place3D: A Full-Cycle LiDAR Placement Pipeline
In this section, we introduce the Surrogate Metric of Semantic Occupancy Grids (M-SOG) to assess the 3D perception performance of sensor configurations (Section 3.2). Leveraging M-SOG, we propose a novel optimization strategy for multi-LiDAR placement (Section 3.3). Our approach is theoretically grounded; we provide an optimality certification to ensure that optimized placements are close to the global optimum (Section 3.4). Figure 1 depicts an overview of our Place3D framework.
3.1 Preliminary Formulation of Surrogate Metric
Region of Interest. Following the literature of sensor placement [34, 57], we define the Region of Interest (ROI) for perception as a finite 3D cuboid space $[l,w,h]$ with ego-vehicle in the center, and divide the ROI space $S$ into voxels with resolution $[\delta_{l},\delta_{w},\delta_{h}]$ as $S=\{v_{1},v_{2},...,v_{N}\}$, where $N=\frac{l}{\delta_{l}}\times\frac{w}{\delta_{w}}\times\frac{h}{\delta_{h}}$ denotes the total number of voxels.
Probabilistic Occupancy Grids (POG). We define POG as the joint probability of all non-zero voxels in the ROI as $p_{POG}=p(v_{1},v_{2},...,v_{N})=\prod_{i=1,\;p(v_{i})\neq 0}^{N}{p(v_{i})}$, where each voxel is assumed to be independently and identically distributed. The voxel $v_{i}$ is occupied if it is within any 3D bounding box $b_{c}$ of object class $c$ at frame $t$, which is denoted as $v_{i}=b_{c}^{(t)}$. Therefore, the probability of being occupied for voxel $v_{i}$ among all $T$ frames can be estimated as $p(v_{i})=\sum_{t=1}^{T}\mathds{1}(v_{i}=b_{c}^{(t)})/{T}$.
Probabilistic Surrogate Metric. Now given some LiDAR placement $L$, the conditional POG can be found as the conditional joint distribution $p_{POG|L}=p(v_{1},v_{2},...,v_{N}|L)$ in the literature [34]. To show how much uncertainty the placement $L$ can reduce in the ROI, the naive surrogate metric S-MIG [34] is introduced as the reduction of the entropy as $-H_{POG|L}=-\sum_{i=1}^{N}H(v_{i}|L)$ with constant $H_{POG}=\sum_{i=1}^{N}H(v_{i})$, where $H(v_{i}),H(v_{i}|L)$ can be found through the entropy of Bernoulli distribution $p(v_{i}),p(v_{i}|L)$, respectively.
@@ 3.2 M-SOG: Surrogate Metric of Semantic Occupancy Grids
The naive surrogate metric S-MIG [34] employs 3D bounding boxes as priors to understand 3D scene distribution. However, this approach exhibits substantial deviations from the actual physical boundaries of the objects and overlooks the occlusion relationships between objects and the environment in LiDAR applications. Furthermore, it is limited to detection tasks only. To overcome these limitations, we propose the Surrogate Metric of Semantic Occupancy Grids to assess and optimize LiDAR placement for 3D scene understanding tasks. Leveraging the Semantic Occupancy Grids derived from diverse environments, our approach can be expanded to tackling adverse conditions.
@@
Semantic Occupancy Grids (SOG). Given a 3D driving scene with a set of semantic classes $Y=\{y_{1},y_{2},...,y_{M}\}$, where $M$ represents the total number of semantic classes in the scene and assuming that each voxel can only be occupied by one semantic tag for each frame, we denote $y^{(t)}(v_{i})$ as the semantic class of voxel $v_{i}\in\{v_{1},v_{2},...,v_{N}\}$ at time frame $t\in\{1,2,...,T\}$. Notably, empty voxels are also considered a semantic class. Accordingly, the set of voxels occupied by semantic class $y_{c}$ at frame $t$ is defined as $S_{y_{c}}^{(t)}=\{v_{i}|y^{(t)}(v_{i})=y_{c}\},\;c=1,2,...,M,\;t=1,2,...,T$. Then, we introduce SOG to describe the total semantic voxel distribution:
$S_{SOG}^{(t)}=\{S_{y_{1}}^{(t)},S_{y_{2}}^{(t)},...,S_{y_{M}}^{(t)}\},\;t=1,2,...,T~{}.$ (1)
@@
Probabilistic Semantic Occupancy Grids. In contrast to the Bernoulli distribution POG [34], we propose a more accurate Multinomial distribution: Probabilistic Semantic Occupancy Grids (P-SOG), denoted as $p_{SOG}$, to represent the probability distribution of voxels belonging to certain semantic classes. Before estimating $p_{SOG}$, we first traverse all frames from given scenes to obtain the probability $\hat{p}$ for each voxel $v_{i}$ occupied by each semantic class $y_{c}$:
$\hat{p}(v_{i}=y_{c})=\frac{|\{t\in\{1,2,\ldots,T\}|v_{i}\in S_{y_{c}}^{(t)}\}|}{T},~{}c=1,2,\ldots,M~{}.$ (2)
@@
Notably, we denote voxel $v_{i}$ being occupied by semantic class $y_{c}$ as $v_{i}=y_{c}$, and $\hat{p}$ indicates estimated distributions from observed samples, whereas $p$ refers to the statistical parameters to be estimated, which are unknown and non-random. We compute the joint probability of all non-zero voxels in the ROI to estimate the $p_{SOG}$. Following the literature [34], the joint distribution over the voxel set $S$ is:
$\hat{p}_{SOG}=\hat{p}(v_{1},v_{2},...,v_{N})=\prod_{i=1,\;\hat{p}(v_{i})\neq 0}^{N}{\hat{p}(v_{i})}~{}.$ (3)
The uncertainty of Probabilistic Semantic Occupancy Grids reflects the 3D scene understanding capability of sensors within a given scene. To quantify this uncertainty, we calculate the entropy for the probability distribution of each voxel $v_{i}$ in $\hat{p}_{SOG}$ as: $\hat{H}(v_{i})=-\sum_{c=1}^{M}\hat{p}(v_{i}=y_{c})\log\hat{p}(v_{i}=y_{c})$.
Based on the independent and identically distributed assumption of SOG, the entropy of P-SOG over the voxel set $S$ is given by: $\hat{H}_{SOG}=\mathds{E}_{v_{i}\sim p_{S}}\sum_{i=1}^{N}\hat{H}(v_{i})$, where $p_{S}$ denotes the $p_{SOG}$ over the voxel set $S$. From the perspective of density estimation, the true P-SOG and its corresponding entropy can be estimated as $p_{SOG}=\hat{p}_{SOG}$ and $H_{SOG}=\hat{H}_{SOG}$, respectively.
@@
Surrogate Metric of Semantic Occupancy Grids. To evaluate LiDAR placements, we further analyze the joint probability distribution of voxels covered by LiDAR rays with varied LiDAR placements. Leveraging the Amanatides and Woo’s Fast Voxel Traversal Algorithm [1], we obtain the voxels covered by LiDAR rays as $S|L_{j}=\text{FastVT}(S,L_{j})=\{v_{1}^{L_{j}},v_{2}^{L_{j}},...,v^{L_{j}}_{N_{j}}\}$ given LiDAR configuration $L=L_{j}$, $j=1,2,...,J$, where $j$ indexes the LiDAR placements, $J$ is the number of total configurations, and $N_{j}$ is the number of voxels covered by rays of LiDAR configuration $L=L_{j}$. Then, the semantic entropy distribution of P-SOG over the voxel set $S|L_{j}$ can be estimated as:
$H_{SOG}^{L_{j}}=\mathds{E}_{v_{i}^{L_{j}}\sim p_{S|L_{j}}}\sum_{i=1}^{N_{j}}\hat{H}(v_{i}^{L_{j}})~{}.$ (4)
@@
We further define the information gain of 3D scene understanding as $\Delta H=H_{SOG}-H_{SOG}^{L_{j}}$ to describe the perception capability of LiDAR configuration $L_{j}$. Since $H_{SOG}$ is a constant given certain 3D semantic scenes with the fixed P-SOG distribution, we propose the normalized Surrogate Metric of Semantic Occupancy Grids (M-SOG) as follows to evaluate the perception capability:
$M_{SOG}(L_{j})$ $=-\frac{1}{N_{j}}H_{SOG}^{L_{j}}=-\frac{1}{N_{j}}\mathds{E}_{v_{i}^{L_{j}}\sim p_{S|L_{j}}}\sum_{i=1}^{N_{j}}H(v_{i}^{L_{j}})$
$=\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\sum_{c=1}^{M}\hat{p}(v_{i}^{L_{j}}=y_{c})\log\hat{p}(v_{i}^{L_{j}}=y_{c})~{}.$ (5)